### v1

# GUT-Scale Smooth Hybrid Inflation with a Stabilized Modulus in Light of ACT and SPT Data

Abstract:
We analyze a generalized framework of smooth F-term hybrid inflation (smFHI) consistent with gauge coupling unification within the Minimal Supersymmetric Standard Model (MSSM). The embedding of the model in two specific Supergravity settings addresses at the same time the $\eta$ problem and the compatibility with the recent ACT or SPT data. The one relies on the choice of a shift-symmetric Kähler potential for the inflaton which revitalizes the SUSY predictions of smFHI, whereas the other employs a Kähler potential associated with an hyperbolic Kähler manifold. An essential role in both our constructions is played by a decoupled superheavy field without superpotential and Kähler potential inspired by string- and D-brane–-based models. Our proposal can be realized for a variety of representations for the Higgs fields involved in smFHI and assures monotonic inflationary potential.
Cosmology of Theories beyond the SM, Supergravity Models
1 Introduction
Among the various versions [1, 2] of F-term hybrid inflation (FHI) the so-called smooth [3, 5, 4, 6, 7, 8, 9, 10, 11, 13, 16, 12, 14, 15] is well distinguishable and observationally more promising – for a variant called pseudosmooth tribrid inflation see Ref. [17, 18]. This is because the inflationary path displays a classical inclination which drives the inflaton $S$ towards its vacuum smoothly – contrary to the standard [19] and shifted [20, 21, 22, 25, 23, 24] versions of FHI which are tied to a classically flat trajectory which develops an instability at some point terminating abruptly FHI. As a consequence, the (scalar) spectral index, $n_{\rm s}$ within smooth FHI (smFHI) in the context of supersymmetry (SUSY) turns out to be significantly lower than the other two models. Namely, we obtain [3, 2]
$n_{\rm s}\simeq 1-5/3N_{\star}=0.967\>\>\mbox{with}\>\>N_{\star}=50$ (1)
being the number of e-foldings that the pivot scale $k_{\star}$ suffered during smFHI. Moreover, smFHI offers a parametric flexibility so that the vacuum expectation values (v.e.vs) of the inflaton-accompanying Higgs fields are less restricted by the normalization of the power spectrum of the curvature perturbation compared to the situation in the standard FHI [2]. As a result, these v.e.vs can become [3, 2] precisely equal to the value suggested by the unification of the gauge coupling constants within SUSY Grand Unified Theories (GUTs) adopting the particle content of the Minimal SUSY Standard Model (MSSM) up to the unification scale.
The result in Eq. (1) is tantalizing close to the central value reported by the Planck data [26] combined with BICEP/Keck (BK) [27] according to which
$n_{\rm s}\simeq 0.9652\pm 0.0084~~\mbox{at 95$\%$ \emph{confidence level} ({\sf\small c.l.})}.$ (2)
This value, however, is substantially lower than the so-called P-ACT-LB-BK18 data which includes the latest Data Release 6 (DR6) from the Atacama Cosmology Telescope (ACT) [28, 29], combined with the cosmic microwave background (CMB) measurements by Planck [26] and BICEP/Keck (BK) [27], together with the Dark Energy Spectroscopic Instrument (DESI) Baryon Acoustic Oscillation (BAO) results [30]. Namely, the P-ACT-LB-BK18 data entails [29]
$n_{\rm s}=0.9743\pm 0.0068\>\Rightarrow\>0.967\lesssim n_{\rm s}\lesssim 0.981\>\>\mbox{at 95$\%$ c.l.}$ (3)
On the other hand, the South Pole Telescope (SPT) announced recently [31] new measurements. Combining these with the ones from Planck and ACT, we obtain the P-ACT-SPT data which dictates
$n_{\rm s}=0.9684\pm 0.006\>\Rightarrow\>0.962\lesssim n_{\rm s}\lesssim 0.974\>\>\mbox{at 95$\%$ c.l.}$ (4)
A wide stream of works [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 50, 51, 52, 53, 54, 55, 56, 57, 58, 60, 49, 59, 61, 62, 63, 64, 65, 66, 67, 68, 69, 32, 33, 34, 35, 36] appeared recently trying to reconcile several well-motivated inflationary models with the observational limits above – mainly that of Eq. (3).
Aspiring to establish realistic versions of smFHI, we have to remind that the inclusion of supergravity (SUGRA) corrections with canonical Kähler potentials – within the so-called minimal SUGRA (mSUGRA) scenario – typically pushes $n_{\rm s}$ above the aforementioned observational margins. This occurs despite the fact that the generation of a mass term for $S$ is avoided thanks to a mutual cancellation. This upward shift in $n_{\rm s}$ can, however, be circumvented if the analysis is restricted to scales well below the SUSY GUT scale. In this regime [16] the P-ACT-LB-BK18 results in Eq. (3) can be marginally reproduced for $N_{\star}=55$. Alternatively, employing a quasi-canonical Kähler potential [11, 6, 13, 12] with a suitable choice of the sign of the next-to-minimal terms allows smFHI to agree with observational data above. This setup, however, leads to hilltop-type [70] solutions over most of the parameter space introducing thereby, an inherent fine-tuning of the initial conditions, related to the maximum – minimum structure of the inflationary potential – cf. Ref. [2, 6]. Insisting on the one step inflationary paradigm – cf. Ref. [71] – a more drastic, and arguably more radical, approach to circumvent the $n_{\rm s}$ problem of smFHI is the inclusion of additional fields into the model which exhibit no-scale-type Kähler potentials. These fields naturally arise as moduli in string theory [72, 73] and can be assumed stabilized [75, 2, 76, 74] during smFHI. They contribute extra terms to the SUGRA corrections of the inflationary potential [78, 79, 77, 2, 75] affecting thereby the inflationary observables. An added advantage of this construction is the possibility of evading the notorious $\eta$ problem [80] of FHI by constraining the coefficients of $K$ to natural values (of order unity).
Working along the lines of the latter idea, we here present novel variants of smFHI which may be reconciled with the data in Eq. (4) or Eq. (3). In the first variant, the inclusion of one modulus together with a shift-symmetric $K$ [81, 82] for the inflaton assist to identify the $n_{\rm s}$ value with that in Eq. (1). We call this setting shift-symmetric SUGRA (shSUGRA). In the second variant, the utilization of an hyperbolic Kähler potential [78, 21, 74, 83] for the inflaton assures a slight positive shift of $n_{\rm s}$ without resorting to hilltop solutions. This scenario is named $N$-depended SUGRA (NSUGRA). Our framework also enables us to extend smFHI to alternative representations of the relevant Higgs fields – cf. [5, 9, 10, 14] – and to explore various choices of the associated exponents – cf. Ref. [7, 16]. Furthermore, we verify the viability of smFHI with a cutoff scale in the superpotential as high as the Planck mass, $m_{\rm P}$. A notable byproduct of our setting is that the v.e.vs of the Higgs fields can be fixed exactly at the SUSY GUT scale. Imposing this condition significantly reduces the parameter space of the model, thereby enhancing its predictive power.
In what follows, we present the salient features of our set-up in Sec. 2 and derive the corresponding inflationary potential (Sec. 3). We then outline the observational constraints relevant to our setup (Sec. 4), before proceeding to our numerical analysis (Sec. 5) and concluding remarks (Sec. 6). In Appendices A and B we propose a sample stabilization of the modulus before smFHI and clarify the status of GUT-scale smFHI within mSUGRA.
2 Set-up
We here present the building blocks of our set-up. Namely, in Sec. 2.1 we introduce the superpotentials for the two types of smFHI whereas in Sec. 2.2 we quote the Kähler potentials which define the two SUGRA versions, shSUGRA and NSUGRA.
2.1 Superpotentials
Smooth FHI can be realized by adopting one of the following superpotentials [3, 7, 16, 14, 10]:
[…]
Here $p$ and $q$ are integers with $p>1$ and $q>2$. In case of smFHI-I, the fields $\bar{\Phi}$ and $\Phi$ form a pair of left-handed superfields belonging to non-trivial conjugate representations of the GUT gauge group $\mathbb{G}$, reducing its rank through their v.e.vs – see e.g. Ref. [5, 6, 15]. By contrast, in smFHI-IIthe Higgs superfield $\Psi=\Psi^{\rm a}{\sf T}^{\rm a}$ transforms in the adjoint representation of $\mathbb{G}$, which may be identified with e.g. $SU(2)_{\rm R}$ of $SU(3)_{\rm c}\times SU(2)_{\rm L}\times SU(2)_{\rm R}\times U(1)_{B-L}$ [9] or $SU(5)$ [23, 10, 14], or $SU(4)_{\rm c}$ of $SU(4)_{\rm c}\times SU(2)_{\rm L}\times SU(2)_{\rm R}$ or $SU(4)_{\rm c}\times SU(2)_{\rm L}\times U(1)_{\rm R}$ [25]. For the adjoint representation, we adopt the normalization
$\mbox{\sf Tr}({\sf T}^{\rm a}{\sf T}^{\rm b})=\delta^{\rm ab}/2~~\mbox{where}~~{\rm a}=1,...,{\cal N}^{2}-1.$ (6)
Here Tr denotes the trace of the Hermitian matrix ${\sf T}^{\rm a}$, and $\rm a$} runs over the dimensionality of the corresponding algebra. Consequently, the kinetic term of $\Psi$ takes the form $\mbox{\sf Tr}(\dot{\Psi})^{2}=(\dot{\Psi}^{\rm a})^{2}/2$, where the dot indicates derivative with respect to cosmic time.
The field $S$ is a $\mathbb{G}$-singlet left-handed superfield, and $M_{*}$ represents an effective cutoff scale, naturally comparable to the string scale $M_{\rm S}=5\times 10^{17}~{\mbox{\rm GeV}}$. We also explore a simplified version of smFHI [13, 14], where
$M_{*}=\begin{cases}\kappa^{1/2(1-p)}m_{\rm P}&\mbox{for smFHI-I}\\ \kappa^{1/(2-q)}m_{\rm P}&\mbox{for smFHI-II}.\end{cases}$ (7)
To ensure the perturbativity we impose on $\kappa$ the condition $\kappa<\sqrt{4\pi}$.
The superpotential in Eq. (5) is consistent with a continuous $R$ symmetry [19], under which $S$ and $W$ carry the same charge, while $(\bar{\Phi}\Phi)^{p}$ or $\Psi^{q}$ remain neutral. In addition, we impose on $W$ an extra global $\mathbb{Z}_{p}$ or $\mathbb{Z}_{q}$ discrete symmetry, under which $\Phi\rightarrow e^{2\pi i/p}\Phi$ or $\Psi\rightarrow e^{2\pi i/q}\Psi$ for smFHI-I or smFHI-II, respectively. As a result, only the $p$ or $q$ powers of $\bar{\Phi}\Phi$ or $\Psi$ correspondingly are allowed.
2.2 Kähler potentials
The full Kähler potentials of our models include separated contributions for the inflaton, $K_{\rm I}$, the stabilized modulus, $^{\hat{K}}$, and the Higgs fields $K_{\rm H}$. It can be written as
$K=K_{\rm I}+{\hat{K}}+K_{\rm H}\>\>\mbox{with}\>\>K_{\rm H}=\begin{cases}|\Phi|^{2}+|\bar{\Phi}|^{2}&\mbox{for smFHI-I}\\ \mbox{\sf Tr}|\Psi|^{2}&\mbox{for smFHI-II.}\end{cases}$ (8)
The form of $K_{\rm I}$ depends on the SUGRA scenario. We consider two specific choices which include just quadratic terms as regards $S$. Namely,
  • For shSUGRA $K_{\rm I}$ assumes the following form
    $K_{\rm I}=-\frac{1}{2}\hat{Z}\left(S-S^{*}\right)^{2},$ (9)
    which breaks the $R$ symmetry of $W$. However, it is invariant under the following symmetries: $S\rightarrow S+c$ and $S\rightarrow-S$, where $c$ is a real number.
  • For NSUGRA $K_{\rm I}$ assumes the following form
    $K_{\rm I}=Nm_{\rm P}^{2}\ln\left(1+{\hat{Z}|S|^{2}}/{Nm_{\rm P}^{2}}\right),$ (10)
    which respects the $R$ symmetry of $W$ and parameterizes either a compact ($N>0$) or non-compact (hyperbolic, $N<0$) moduli Kähler manifold.
The functions $^{\hat{K}}$ in Eq. (8) and $\hat{Z}$ in Eqs. (9) and (10) above encode the dependence of $K$ on the stabilized modulus $h$. We assume that $h$ is charged under some extra symmetry and so it is not present in $W$ [78, 79]. Motivated by superstring and D-brane models [72, 73], we adopt the ansatz
$\hat{Z}=\left({f(h)+f^{*}(h^{*})}\right)^{\alpha}\>\>\mbox{and}\>\>{\hat{K}}=\beta m_{\rm P}^{2}\ln\left(f(h)+f^{*}(h^{*})\right)\>\>\mbox{with}\>\>\beta<0,$ (11)
where $f$ is an analytic function of $h$, $\beta$ is taken to be an integer and $\alpha$ a rational number. The restriction $\beta<0$ ensures positivity of the kinetic terms of $h$. $^{\hat{K}}$ parameterizes the Kähler manifold $SU(1,1)/U(1)$ with constant scalar curvature $2/\beta$. We assume that $h$ is stabilized during smFHI through some mechanism – see Appendix A for an example. In particular, we set
$\langle{f(h)}\rangle_{\rm I}=\langle{f^{*}(h^{*})}\rangle_{\rm I}=1/2~~\mbox{so as}~~\langle{\hat{Z}}\rangle_{\rm I}=1~~\mbox{and}~~\langle{{\hat{K}}}\rangle_{\rm I}=0,$ (12)
where the symbol $\langle{Q}\rangle_{\rm I}$ means the value of the quantity $Q$ during smFHI.
Taking into account the form of the various $K$’s above, we can define the canonically normalized real fields $\sigma,\bar{\sigma},\phi$ and $\psi$ as follows
$S=(\sigma+i\bar{\sigma})/\sqrt{2}~~\mbox{and}~~\begin{cases}\Phi=\bar{\Phi}=\phi/2&\mbox{for smFHI-I,}\\ \Psi^{1}=\psi&\mbox{for smFHI-II,}\end{cases}$ (13)
where the MSSM-neutral component of $\Psi$ is aligned along the ${\rm a}=1$ direction. The imaginary part of $S$ can be absorbed through an $R$-transformation in case of NSUGRA. However, $R$-symmetry is explicitly broken within shSUGRA – see Eq. (9) – and so the stability of the direction $\bar{\sigma}=0$ has to be checked as we see in Sec. 3.1 below. Since $h$ is stabilized during smFHI there is no hSh-S mixing within NSUGRA [75, 74, 2].
3 Inflationary Potential
We here introduce the SUGRA potential of our models in Sec. 3.1 and pay special attention in computing the SUSY contribution in Sec. 3.2. Finally, we obtain a compact expression for the potential in Sec. 3.3 which supports the various versions of smFHI discussed in our work.
3.1 SUGRA Potential
The F-term SUGRA scalar potential is given by
$V_{\rm SUGRA}=e^{K/m_{\rm P}^{2}}\left(K^{I\bar{J}}{\rm F}_{I}{\rm F}^{*}_{\bar{J}}-3|W|^{2}/m_{\rm P}^{2}\right)+\frac{g^{2}}{2}\mbox{$\sum_{a}$}{\rm D}_{a}{\rm D}_{a}$ (14)
where the subscript $I~[\bar{J}]$ denotes derivation with respect to (w.r.t) the complex scalar field $X_{I}~[X_{\bar{J}}^{*}]$ which corresponds to the chiral superfield $X_{I}$ with $X_{I}=S,h$ and $\Phi,\bar{\Phi}$ or $\Psi$. The matrix $K^{I\bar{J}}$ is the inverse of the Kähler metric $K_{I\bar{J}}=K_{,X^{I}X^{\bar{J}}}$ – with the symbol $X^{J}$ as subscript denoting derivation with respect to (w.r.t) $X^{J}$ –, i.e., $K^{I\bar{J}}K_{I\bar{L}}=\delta^{\bar{J}}_{\bar{L}}$. Also the F and D terms are
${\rm F}_{I}=W_{,X^{I}}+K_{,X^{I}}W/m_{\rm P}^{2}\>\>\mbox{and}\>\>{\rm D}_{a}=X^{I}\left(T_{a}\right)^{J}_{I}K_{J},$ (15)
with $K_{J}={K_{,X^{J}}}$. Here, the index $I$ should not to be confused with the subscript “I”, which is used to denote evaluation during smFHI.
smFHI takes place along a D-flat inflationary trajectory defined by the conditions
$\begin{cases}\langle{|\Phi|-|\bar{\Phi}|}\rangle_{\rm I}=0&\mbox{for smFHI-I,}\\ \langle{|\Psi^{\dagger}\mbox{\sf Tr}[\Psi,{\sf T}^{\rm a}]|}\rangle_{\rm I}^{2}=0&\mbox{for smFHI-II,}\end{cases}$ (16)
which remain unaffected by the SUGRA corrections, as we consider canonical kinetic terms for the Higgs fields in Eq. (8). For smFHI-II the condition is automatically satisfied, owing to the antisymmetric nature of the $SU({\cal N})$ structure constants and the fact that all component fields which are not singlets under the Standard Model gauge group vanish during smFHI.
Taking into account that $W_{,h}=0$, according to our set-up in Sec. 2.1, and that $\Phi,\bar{\Phi}\ll S$ for smFHI-I and $\Psi\ll S$ for smFHI-II, $V_{\rm F}$, in Eq. (14) can be written as
$V_{\rm F}\simeq V_{\rm F0}$ $+$ $K^{SS^{*}}\left(|WK_{S}/m_{\rm P}^{2}|^{2}+W_{,S}W^{*}K_{S^{*}}/m_{\rm P}^{2}+{\rm c.c}\right)$ (17)
$+$ $K^{Sh^{*}}WK_{h}/m_{\rm P}^{2}\left(W_{,S^{*}}+W^{*}K_{S^{*}}/m_{\rm P}^{2}\right)$
$+$ $K^{hS^{*}}WK_{h^{*}}/m_{\rm P}^{2}\left(W_{,S}+WK_{S}/m_{\rm P}^{2}\right)$
$+$ $K^{hh^{*}}|W|^{2}K_{h}K_{h^{*}}/m_{\rm P}^{4}-3|W|^{2}/m_{\rm P}^{2},$
where the $V_{\rm F0}$ is the SUSY part of $V_{\rm F}$ which is given in Sec. 3.2. The remaining part of $V_{\rm F}$ dependents on the SUGRA scenario. Focusing on the $(S,h)$ subspace of our field space we below derive $\langle{K_{I\bar{J}}}\rangle_{\rm I}$ and $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$. Namely,
  • For shSUGRA we find
    $\langle{K_{I\bar{J}}}\rangle_{\rm I}={\sf diag}(1,-\beta\langle{|f_{,h}|}\rangle_{\rm I}^{2})~~\mbox{and}~~\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}=-(3+\beta)M^{4}\sigma^{2}/2m_{\rm P}^{2},$ (18)
    where we take into account $\langle{K_{S}}\rangle_{\rm I}=\langle{K_{Sh}}\rangle_{\rm I}=0$ and $\langle{K_{SS}}\rangle_{\rm I}=1$. Note that $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$ is independent of $\langle{f_{,h}}\rangle_{\rm I}$ although this is present in $\langle{K_{I\bar{J}}}\rangle_{\rm I}$. It also exhibits a runaway behavior for $\beta=0$ as first noticed in Ref. [82]. From the complete version of $V_{\rm F}$ we can also derive the effective mass of $\bar{\sigma}$ in Eq. (13) along the direction $\langle{\bar{\sigma}}\rangle_{\rm I}=0$. We find
    $m^{2}_{\bar{\sigma}}=(M^{4}/m_{\rm P}^{2})\left(3+(2\alpha-\beta)^{2}/\beta+(\alpha(3+\alpha)-\beta-1)(\sigma/m_{\rm P})^{2}\right),$ (19)
    which depends on $\alpha$ contrary to $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$ in Eq. (18).
  • For NSUGRA we find the following elements of the Kähler metric
    $\langle{K_{SS^{*}}}\rangle_{\rm I}=\left(1+\sigma^{2}/2Nm_{\rm P}^{2}\right)^{-2}\simeq 1,$ (20)
    $\langle{K_{Sh^{*}}}\rangle_{\rm I}=\langle{K_{hS^{*}}}\rangle_{\rm I}=(\sigma/\sqrt{2})\alpha\langle{f_{,h}}\rangle_{\rm I}\left(1+\sigma^{2}/2Nm_{\rm P}^{2}\right)^{-2}\ll 1,$
    $\langle{K_{hh^{*}}}\rangle_{\rm I}=\left(\alpha\sigma^{2}\left(2(\alpha-1)N-{\sigma}^{2}/{m_{\rm P}^{2}}\right)/4N\left(1+\sigma^{2}/2Nm_{\rm P}^{2}\right)^{2}-\beta m_{\rm P}^{2}\right)|\langle{f_{,h}}\rangle_{\rm I}|^{2}$
    $\hskip 36.98866pt\simeq-\beta|\langle{f_{,h}}\rangle_{\rm I}|^{2}m_{\rm P}^{2}.$
    On the other hand, $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$ takes the form
    $\langle{V_{\rm F}-V_{\rm F0}}\rangle_{\rm I}$ $=$ $M^{4}\left(1+\frac{\sigma^{2}}{2Nm_{\rm P}^{2}}\right)^{N}\left(8\beta N^{3}-4N^{2}((\beta-\alpha)(1-\alpha+\beta)N-3\beta)(\sigma/m_{\rm P})^{2}\right$. (21)
    $+$ $2N\left(\beta(3+N+2\alpha N+N^{2})-\alpha N(N+\alpha-2)-\beta^{2}N\right)(\sigma/m_{\rm P})^{4}$
    $+$ $\left.(1+N)^{2}(\beta+\alpha N)(\sigma/m_{\rm P})^{6}\right)/4N^{2}(2\beta N+(\beta+\alpha N)(\sigma/m_{\rm P})^{2}),$
    which is still independent from $\langle{f_{,h}}\rangle_{\rm I}$ similarly to our result in Eq. (18).
3.2 SUSY Contribution
The SUSY F-term scalar potential, $V_{\rm F0}$, which appears in Eq. (17), takes the form
$V_{\rm F0}=K^{SS^{*}}|W_{,S}|^{2}+\begin{cases}K^{\Phi\Phi^{*}}|W_{,\Phi}|^{2}+K^{\bar{\Phi}\bar{\Phi}^{*}}|W_{,\bar{\Phi}}|^{2}~~&\mbox{for smFHI-I,}\\ K^{\psi\psi^{*}}|W_{,\psi}|^{2}~~&\mbox{for smFHI-II,}\end{cases}$ (22)
where $K^{\bar{\Phi}\bar{\Phi}^{*}}=K^{\Phi\Phi^{*}}=K^{\psi\psi^{*}}=1$ according to Eq. (8). Since $\langle{S}\rangle_{\rm I}\ll m_{\rm P}$ we can also set $K^{SS^{*}}\simeq 1$. Taking into account the form of $W$ in Eq. (5), we find
$V_{\rm F0}=\begin{cases}4^{1-2p}p^{2}\sigma^{2}\phi^{2(2p-1)}/M_{*}^{4(p-1)}+\left(M^{2}-4^{-p}\phi^{2p}/M_{*}^{2(p-1)}\right)^{2}&\mbox{for smFHI-I,}\\ d_{q}^{2}M_{*}^{2(2-q)}q^{2}\sigma^{2}\psi^{2(q-1)}/2+\left(M^{2}-d_{q}M_{*}^{2-q}\psi^{q}\right)^{2}&\mbox{for smFHI-II.}\end{cases}$ (23)
whose minimization yields the SUSY vacua
$\langle{\sigma}\rangle=0~~\mbox{and}~~\begin{cases}\langle{\phi}\rangle=2\left(MM_{*}^{(p-1)}\right)^{1/p}&\mbox{for smFHI-I,}\\ \langle{\psi}\rangle=\left(M^{2}M_{*}^{(q-2)}/d_{q}\right)^{1/q}&\mbox{for smFHI-II,}\end{cases}$ (24)
with $d_{q}=|\mbox{\sf Tr}[T_{1}^{q}]|$ serving as a normalization factor that depends on the representation of $\Psi$. At these vacua, the mass of $\sigma$ is found to be
$m_{\rm I}=\left\langle V_{\rm F0,\sigma\sigma}\right\rangle^{1/2}=M^{2}\cdot\begin{cases}\sqrt{2}p\langle{\Phi}\rangle^{-1}&\mbox{for smFHI-I,}\\ q\langle{\Psi}\rangle^{-1}&\mbox{for smFHI-II.}\end{cases}$ (25)
Upon normalizing the fields w.r.t $\langle{\phi}\rangle$ and $\langle{\psi}\rangle$, Eq. (17) can be reformulated as follows:
$V_{\rm F0}=M^{4}\begin{cases}\left(1-({\phi}/{\langle{\phi}\rangle})^{2p})^{2}+4p^{2}(\sigma/\langle{\phi}\rangle)^{2}(\phi/\langle{\phi}\rangle)^{2(2p-1)}\right)&\mbox{for smFHI-I,}\\ \left(1-({\psi}/{\langle{\psi}\rangle})^{q})^{2}+q^{2}(\sigma/\langle{\psi}\rangle)^{2}(\psi/\langle{\psi}\rangle)^{2(q-1)}\right)&\mbox{for smFHI-II.}\end{cases}$ (26)
Inspection of $V_{\rm F0}$ shows that it features at least one inflationary valley (with nonzero $\phi$ or $\psi$) smoothly connected to the SUSY vacuum. Already at tree level, these valleys display a built-in slope that drives $\sigma$ towards the vacuum. Consequently, unlike in standard [19] or shifted [20, 21, 22] FHI, radiative corrections are not required to provide the necessary slope, though they may still contribute subdominantly. Let us, finally, note that no problematic domain walls arise from the spontaneous breaking of the global $\mathbb{Z}_{p}$ $\mathbb{Z}_{q}$] symmetry in Eq. (24), since this breaking occurs already during smFHI.
From here, we proceed with the analysis of the two cases, smFHI-I and smFHI-II, separately.
(a) Type I smFHI
Taking the derivative of $V_{\rm F0}$ w.r.t $\phi$, we obtain
$V_{\rm F0,\phi}=4^{1-2p}M_{*}^{2(1-2p)}p\phi^{2p-3}\left(M_{*}^{2}\phi^{2p}\left(2p(2p-1)\sigma^{2}+\phi^{2}\right)-4^{p}M^{2}M_{*}^{2p}\phi^{2}\right).$ (27)
Assuming $\phi\ll\sigma$ during smFHI-I, the $\phi^{2}$ term in the first parenthesis can be neglected. This leads to
$\langle{V_{\rm F0,\phi}}\rangle_{\rm I}=0~~\Rightarrow~~\langle{\phi}\rangle_{\rm I}=c_{1p}\left(\sigma/\langle{\phi}\rangle\right)^{1/(1-p)}\langle{\phi}\rangle~~\mbox{with}~~c_{1p}=(2p(2p-1))^{(1/2(1-p))}$ (28)
and $\langle{V_{\rm F0,\phi\phi}}\rangle_{\rm I}>0$. Substituting Eq. (28) into Eq. (17), we obtain
$\langle{V_{\rm F0}}\rangle_{\rm I}$ $=$ $M^{4}\left(1-(p-1)c_{2p}(\sigma/\langle{\phi}\rangle)^{-2p/(p-1)}+c_{4p}(\sigma/\langle{\phi}\rangle)^{-4p/(p-1)}\right),$ (29)
$\simeq$ $M^{4}\left(1-(p-1)c_{2p}(\sigma/\langle{\phi}\rangle)^{-2p/(p-1)}\right),$
where the $p$-dependent coefficients are defined as
$c_{2p}=\left(2p^{p}(2p-1)^{2p-1}\right)^{1/(1-p)}~~\mbox{with}~~c_{4p}=c_{1p}^{4p}.$ (30)
(b) Type II smFHI
Following analogous steps, we find
$\langle{V_{\rm F0,\psi}}\rangle_{\rm I}=0~~\Rightarrow~~\langle{\psi}\rangle_{\rm I}=c_{1q}\left(\sigma/\langle{\psi}\rangle\right)^{2/(2-q)}\langle{\psi}\rangle~~\mbox{with}~~c_{1q}=(q(q-1))^{1/(2-q)}.$ (31)
Substituting Eq. (31) into Eq. (17) yields
$\langle{V_{\rm F0}}\rangle_{\rm I}$ $=$ $M^{4}\left(1-(q-2)c_{2q}(\sigma/\langle{\psi}\rangle)^{-2q/(q-2)}+c_{4q}(\sigma/\langle{\psi}\rangle)^{-4q/(q-2)}\right)$ (32)
$\simeq$ $M^{4}\left(1-(q-2)c_{2q}(\sigma/\langle{\psi}\rangle)^{-2q/(q-2)}\right),$
with the $q$-dependent coefficients defined as follows
$c_{2q}=\left(q^{q}(q-1)^{2(q-1)}\right)^{1/(2-q)}~~\mbox{with}~~c_{4q}=c_{1q}^{2q}.$ (33)
3.3 Final Form
The general form of the potential that can drive the different versions of smFHI can be written as
$V_{\rm I}=M^{4}\,\left(\langle{V_{\rm F0}}\rangle_{\rm I}/M^{4}-{c_{2K}}\frac{\sigma^{2}}{2m_{\rm P}^{2}}+{c_{4K}}\frac{\sigma^{4}}{4m_{\rm P}^{4}}-{c_{6K}}\frac{\sigma^{6}}{8m_{\rm P}^{6}}+{c_{8K}}\frac{\sigma^{8}}{16m_{\rm P}^{8}}\right),$ (34)
where $\langle{V_{\rm F0}}\rangle_{\rm I}$ is given by Eqs. (29) and (32) for smFHI-I and smFHI-II respectively, whereas the coefficients $c_{iK}$ with $i=2,4,6$ and $8$ depend on the SUGRA set-up in Eqs. (9) and (10). Expanding the expressions of $V_{\rm F}$ in Eqs. (18) and (21) as series in powers of $\sigma/m_{\rm P}$ we can determine $c_{iK}$ as follows:
  • For shSUGRA we find
    $c_{2K}=3+\beta~~\mbox{and}~~c_{4K}=c_{6K}=c_{8K}=0.$ (35a)
    Given that $\beta$ is negative integer, this scenario is activated only if we set
    $c_{2K}=0~~\Rightarrow~~\beta=-3$ (35b)
    which totally eliminates, the quadratic SUGRA contribution to $\eta$ parameter. On the other hand, thanks to the shift symmetry in Eq. (9) all remaining SUGRA corrections in Eq. (34) vanish and therefore the SUSY results on the inflationary observables reveal. The selected $\beta$ value renders the path $\langle{\bar{\sigma}}\rangle_{\rm I}=0$ well stabilized during smFHI since from Eq. (19) with $\alpha=0$ (for simplicity) we obtain
    $m^{2}_{\bar{\sigma}}=6\left(3+(\sigma/m_{\rm P})^{2}\right)H_{\rm I0}^{2}\gg H_{\rm I0}^{2}~~\mbox{with}~~H_{\rm I0}\simeq M^{2}/\sqrt{3}m_{\rm P},$ (36)
    being the Hubble parameter during smFHI. Moreover, $\bar{\sigma}$ is sufficiently heavy and so it does not interfere with smFHI, ensuring the consistency of our one-field inflationary scenario – see below.
  • For NSUGRA we find
    $c_{2K}$ $=$ ${(\alpha-\beta)^{2}\over\beta}-{2\over N},$ (37a)
    $c_{4K}$ $=$ ${(\alpha-\beta)^{3}\over\beta^{2}}+{7\over 2N}+{1\over N^{2}}+{1\over 2},$ (37b)
    $c_{6K}$ $=$ $\left({\alpha^{4}\over\beta^{3}}-{3\alpha^{3}\over\beta^{2}}+{7\alpha^{2}\over 2\beta}-2\alpha+{\beta\over 2}-{2\over 3}\right)$ (37c)
    $+{1\over N}\left({\alpha^{3}\over\beta^{2}}-{5\alpha^{2}\over 2\beta}+2\alpha-{\beta\over 2}-3\right)-{1\over 3N^{2}},$
    $c_{8K}$ $=$ ${1\over N}\left({\alpha^{5}\over\beta^{4}}-{3\alpha^{4}\over\beta^{3}}+{7\alpha^{3}\over 2\beta^{2}}-{13\alpha^{2}\over 6\beta}+{5\alpha\over 6}-{\beta\over 6}+{3\over 8}\right)$ (37d)
    $+{1\over N^{2}}\left(+{2\alpha^{4}\over\beta^{3}}-{11\alpha^{3}\over 2\beta^{2}}+{11\alpha^{2}\over(2\beta}-{5\alpha\over 2}+{\beta\over 2}+{13\over 12}\right)$
    $+{1\over N^{3}}\left({\alpha^{3}\over\beta^{2}}-{7\alpha^{2}\over 3\beta}+{5\alpha\over 3}-{\beta\over 3}-{11\over 8}\right)-{1\over 12N^{4}}.$
    From Eq. (37a), we see that $c_{2K}$ can be set to zero by constraining $N$ as a function of $\alpha$ and $\beta$, namely
    $c_{2K}=0~~\Rightarrow~~N=N_{0}~~\mbox{with}~~N_{0}={2\beta\over(\alpha-\beta)^{2}}.$ (38)
    Therefore, we are left with only two free parameters $\alpha$ and $\beta$.
Both conditions in Eqs. (35b) and (38) consist an elegant resolution to the infamous $\eta$ problem of FHI, since for $\alpha$ and $\beta$ of order unity – as we show in Sec. 5 – we can totally avoid the relevant term.
4 Constraining the Inflationary Dynamics
The parameters of our models can be constrained by both observational data and theoretical consistency conditions, as described in Secs. 4.1 and 4.2.
4.1 Observational Constraints
Our models of smFHI is viable only if it satisfies a number of observational requirements. Namely:
The number of e-foldings underwent by the pivot scale $k_{\star}=0.05/{\rm Mpc}$ during smFHI, must be sufficient to address the standard problems of the Big Bang cosmology [2, 26]. Specifically,
$N_{\star}=\>-\frac{1}{m^{2}_{\rm P}}\;\int_{\sigma_{\star}}^{\sigma_{\rm f}}\,d\sigma\>\frac{V_{\rm I}}{V_{\rm I,\sigma}}\simeq 61.3+\frac{1-3w_{\rm rh}}{12(1+w_{\rm rh})}\ln\frac{\pi^{2}g_{\rm rh*}T_{\rm rh}^{4}}{30V_{\rm I}(\sigma_{\rm f})}+\frac{1}{4}\ln{V_{\rm I}(\sigma_{\star})^{2}\over g_{\rm rh*}^{1/3}V_{\rm I}(\sigma_{\rm f})},$ (39)
where $\sigma_{\star}$ is the field value when $k_{\star}$ exits the horizon, and $\sigma_{\rm f}$ marks the end of smFHI. Unlike in standard [19] or (semi)shifted FHI [20, 21, 22, 25], smFHI ends smoothly at $\sigma=\sigma_{\rm f}$, when the slow-roll conditions break down. Concretely, $\sigma_{\rm f}$ is determined by
${\sf\footnotesize max}\{\epsilon(\sigma_{\rm f}),|\eta(\sigma_{\rm f})|\}=1,~~\mbox{where}~~\epsilon\simeq{m^{2}_{\rm P}\over 2}\left(\frac{V_{\rm I,\sigma}}{V_{\rm I}}\right)^{2}~~\mbox{and}~~\eta\simeq m^{2}_{\rm P}~\frac{V_{\rm I,\sigma\sigma}}{V_{\rm I}}\cdot$ (40)
In the last part of Eq. (39) we assume that smFHI is followed in turn by an oscillatory phase (due to $S$ field) with mean equation-of-state parameter $w_{S\rm rh}$ radiation and matter domination, $T_{\rm rh}$ is the reheat temperature after smFHI. Given that $w_{\rm rh}$ is found by the standard formula $w_{\rm rh}=(n-2)/(n+2)$ for a power-law potential $\varphi^{n}$, we take for our numerics $w_{S\rm rh}=0$ which corresponds precisely to $n=2$, since $V_{\rm F0}$ in Eq. (23) displays a quadratic dependence on $S$. On the other hand, the oscillations of the Higgs fields give rise to a stiff component of the total matter energy-density – with $w_{\rm Hrh}>1/3$ since $n>4$ – which is overshadowed by that of $S$ [84]. Here it is expected that the inflationary energy density is equally distributed between the oscillatory subsystems of $S$ and the Higgs fields. Motivated by implementations [5, 9, 15] of non-thermal leptogenesis, which may follow smFHI, we set $T_{\rm rh}\simeq(10^{8}-10^{9})~{\mbox{\rm GeV}}$. Although not crucial for the resulting magnitude of $N_{\star}$ we mention that we take for the energy-density effective number of degrees of freedom $g_{\rm rh*}=228.75$ inspired by the MSSM spectrum. With these choices, we obtain $N_{\rm I\star}\simeq 48.2-49.5$.
The amplitude $A_{\rm s}$ of the curvature perturbation power spectrum generated by $\sigma$ during smFHI must be consistent with CMB observations [29], i.e.
$A_{\rm s}=\frac{1}{12\,\pi^{2}m^{6}_{\rm P}}\;\left.\frac{V_{\rm I}^{3}(\sigma_{\star})}{|V^{\prime}_{\rm I}(\sigma_{\star})|^{2}}\right.\simeq 2.1326\cdot 10^{-9}.$ (41)
The remaining key observables are the scalar spectral index $n_{\rm s}$, its running $a_{\rm s}$, and the tensor-to-scalar ratio $r$. These are obtained from the standard expressions
$n_{\rm s}=1-6\epsilon_{\star}\ +\ 2\eta_{\star},\,a_{\rm s}={2}\left(4\eta_{\star}^{2}-(n_{\rm s}-1)^{2}\right)/3-2\xi_{\star}~~\mbox{and}~~r=16\epsilon_{\star},$ (42)
where $\xi\simeq m_{\rm P}^{4}(V^{\prime}_{\rm I}V^{\prime\prime\prime}_{\rm I})/V_{\rm I}^{2}$, and all quantities with a subscript $\ast$ are evaluated at $\sigma=\sigma_{\ast}$. As regards $n_{\rm s}$ we consider implications to our models taking into account separately Eqs. (3) and (4). The remaining observables must lie within the following approximate ranges [29]:
$a_{\rm s}=0.0062\pm 0.0104~~\Rightarrow~~-0.004\lesssim a_{\rm s}\lesssim 0.017~~\mbox{and}~~r\lesssim 0.038,$ (43)
at 95% c.l. Note that both observables above are compatible with their values in our context, as we show in Sec. 5.
4.2 Theoretical Considerations
From the theoretical perspective, the viability of our model can be further refined by imposing the following requirements:
Unification Constraint. If the gauge group $\mathbb{G}$ contains non-Abelian factors beyond those of MSSM, the mass $M_{A}$ of the lightest gauge boson $A_{\mu}$ in the SUSY vacuum – see Eq. (24) – must be consistent with gauge coupling unification within MSSM, namely
$M_{A}\simeq 2\cdot 10^{16}~{\mbox{\rm GeV}}\>\mbox{where}\>M_{A}=\begin{cases}{g\langle{\Phi}\rangle}&\mbox{for smFHI-I,}\\ {gn_{\rm G}\langle{\Psi}\rangle}&\mbox{for smFHI-II,}\end{cases}$ (44)
where $g\simeq 0.7$ is the unified gauge coupling constant, and $n_{\rm G}$ is a group-theoretic normalization factor depending on $\mathbb{G}$. For example, in the case $\mathbb{G}=SU(5)$, one finds $n_{\rm G}=\sqrt{5/6}$.
Boundedness of $V_{\rm I}$. The inflationary potential must be bounded from below to avoid the possibility of a disastrous runaway of the system to infinite values of the inflaton field. This requirement also facilitates the possibility that the system may eventually undergo an inflationary expansion under generic initial condition.
Convergence of $V_{\rm I}$. The expansion of $V_{\rm I}$ in Eq. (34) should converge at least up to the relevant field values $\sigma\sim\sigma_{\ast}$. This is guaranteed if, for such values, each successive term in the aforementioned expansion is smaller than the preceding one. In practice, this requirement can be satisfied if our results remain immune from the contributions with coefficients $c_{iK}$ with $i>6$.
Monotonicity of $V_{\rm I}$. Depending on the values of the coefficients $c_{iK}$ in Eq. (34), the potential $V_{\rm I}$ may either remain monotonic in $\sigma$ or develop a local minimum and maximum. In the scenarios of interest here, $V_{\rm I}$ turns out to be monotonic, avoiding the complications of metastable extrema – see Sec. 5.3.
5 Results
The discussed scenarios of smFHI depend on the following parameters
$M,~M_{*}~~\mbox{and}~~p~[q]~~\mbox{for smFHI-I\ [smFHI-II] and}~~\alpha,~\beta~~\mbox{only for NSUGRA.}$
Recall that the resolution of $\eta$ problem forces us to determine $\beta$ from Eq. (35b) for shSUGRA and $N$ as a function of $\alpha$ and $\beta$ via Eq. (38) for NSUGRA. The unification condition of Eq. (44) in conjunction with Eq. (24) allows us to obtain $M_{*}$ in terms of $M$ as follows
$M_{*}=\begin{cases}\left(\langle{\Phi}\rangle^{p}/M\right)^{1/(p-1)}&\mbox{for smFHI-I,}\\ \left(\langle{\Psi}\rangle d_{q}/n_{\rm G}M^{2/p}\right)^{p/(p-2)}&\mbox{for smFHI-II,}\end{cases}$ (45)
where $d_{q}=2^{-q}15^{-q/2}\left(3\,2^{q}(-M)^{q}+2\,3^{q}M^{q}\right)$ – note that for smFHI-II we confine ourselves in the case with $\mathbb{G}=SU(5)$. Moreover, $M$ can be determined by enforcing Eq. (41) – note that Eq. (39) constrains $\sigma_{\star}$ which is an internal parameter of the inflationary setting. Therefore, the number of the free parameters of our models can be reduced by two.
5.1 shSUGRA
In the context of shSUGRA $V_{\rm I}$ is given by the general form of Eq. (34) with imposition of the conditions in Eqs. (35a) and (35b). As a consequence, $V_{\rm I}$ coincides with $\langle{V_{\rm F0}}\rangle_{\rm I}$ in Eq. (29) or (32) for smFHI-I or smFHI-II respectively and so the observables are totally identical to those obtained within rigid SUSY, i.e., they only depend on the parameters $p$ or $q$ for smFHI-I or smFHI-II respectively. Our results are summarized for each type of smFHI separately below.
(a) Type I smFHI
In this case the slow-roll parameters can be found by inserting $V_{\rm I}$ as described above into Eq. (40) with result
$\epsilon_{0}=2(2p)^{2}c_{2p}^{2}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\phi}\rangle}\right)^{4p/(1-p)}~~\mbox{and}~~\eta_{0}=\frac{4c_{2p}(1-3p)p}{(p-1)}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\phi}\rangle}\right)^{2p/(1-p)}.$ (46)
The number of e-foldings $N_{\star}$ is found by applying the left-most expression in Eq. (39) as follows
$N_{\star}=\frac{(p-1)\sigma_{\star}^{4+2/(p-1)}\langle{\phi}\rangle^{2p/(1-p)}}{8c_{2p}m_{\rm P}^{2}p(2p-1)}~~\Rightarrow~~\sigma_{\star}=2^{\frac{3(p-1)}{4p-2}}\left(\frac{(p-1)\langle{\phi}\rangle^{-2p/(p-1)}}{c_{2p}m_{\rm P}^{2}N_{\star}p(2p-1)}\right)^{\frac{1-p}{4p-2}}.$ (47)
Solving w.r.t $M$ Eq. (41) we find $M$ as follows
$M\simeq 2m_{\rm P}^{3/2}\sqrt{\pi pc_{2p}\sqrt{3A_{\rm s}}}\sigma_{\star}^{-1/2}\left(\langle{\phi}\rangle/\sigma_{\star}\right)^{p/(p-1)},$ (48)
where $\sigma_{\star}$ is given by Eq. (47) – its explicit replacement in the expression above yields a rather lengthly final result. Substituting also Eq. (47) into Eq. (42) we find the predictions for $n_{\rm s}$ and $a_{\rm s}$ which become
$n_{\rm s0}=1-{3p-1\over 2p-1}\frac{1}{N_{\star}}~~\mbox{and}~~a_{\rm s0}=-{3p-1\over 2p-1}\frac{1}{N_{\star}^{2}}$ (49)
whereas $r$ remains rather suppressed. The expression of $n_{\rm s}$ is in agreement with the one found in Ref. [16]. Numerical and more precise values for the quantities above are accumulated in Table 1 for $p=2,3$ and $4$. We observe that $\sigma_{\star},\sigma_{\rm f}$ and $M_{*}$ are of order $M_{\rm S}$ whereas $M$ and $m_{\rm I}$ are of order $1~{\mbox{\rm YeV}}$ and $0.1~{\mbox{\rm YeV}}$ increasing slightly with $p$ – recall that $1~{\mbox{\rm YeV}}=10^{15}~{\mbox{\rm GeV}}$. The $n_{\rm s}$ values are impressively consistent with the P-ACT-SPT data in Eq. (4) and marginally disfavored by the P-ACT-LB-BK18 data in Eq. (3) independently from $p$. On the other hand, $a_{\rm s}$ and $r$ safely satisfy the observational bounds in Eq. (43). Lastly, the identification of $M_{*}$ with $m_{\rm P}$ is not possible since $\kappa$ defined in Eq. (7) violates the perturbative limit.
Cases Type I smFHI Type II smFHI
$p=2$ $p=3$ $p=4$ $q=3$ $q=5$ $q=7$
$\sigma_{\star}/10^{17}~{\mbox{\rm GeV}}$ $2.65$ $3.32$ $3.47$ $1.49$ $2.64$ $2.95$
$M_{*}/M_{\rm S}$ $1.57$ $0.27$ $0.16$ $8.5$ $0.14$ $0.07$
$\kappa$ $\sqrt{4\pi}$ $0.57$ $\sqrt{4\pi}$
$M/10^{15}~{\mbox{\rm GeV}}$ 11 $1.3$ $1.35$ $0.7$ $1.1$ $1.2$
$\sigma_{\rm f}/10^{17}~{\mbox{\rm GeV}}$ $1.34$ $1.45$ $1.42$ $0.85$ $1.22$ $1.24$
$N_{\rm I*}$ $48.8$ $49.3$ $49$ $48.6$ $48.9$ $48.9$
$n_{\rm s}$ $0.966$ $0.968$ $0.968$ $0.964$ $0.967$ $0.968$
$\alpha_{\rm s}/10^{-4}$ $6.7$ $6.4$ $6.3$ $7.4$ $6.7$ $6.4$
$r/10^{-6}$ $1.1$ $2.4$ $3.8$ $0.197$ $1.3$ $2.1$
$m_{\rm I}/10^{14}~{\mbox{\rm GeV}}$ $1.1$ $2.4$ $3.6$ $0.45$ $1.94$ $3.4$
Table 1: Input and output parameters compatible with Eqs. (39), (41) and (44), within SUSY or shSUGRA for smFHI-I or smFHI-II and selected $p$ or $q$ values respectively.
(b) Type II smFHI
For smFHI-II, a similar analysis yields for the slow-roll parameters in Eq. (40)
$\epsilon_{0}=2q^{2}c_{2q}^{2}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\psi}\rangle}\right)^{4q/(2-q)}~~\mbox{and}~~\eta_{0}=\frac{2c_{2q}(2-3q)q}{(q-2)}\left(\frac{m_{\rm P}}{\sigma}\right)^{2}\left(\frac{\sigma}{\langle{\psi}\rangle}\right)^{2q/(2-q)},$ (50)
while $\sigma_{\star}$ as a function of $N_{\star}$ can be found as follows
$N_{\star}\simeq\frac{(q-2)\sigma_{\star}^{4+4/(q-2)}\langle{\psi}\rangle^{2q/(2-q)}}{8c_{2q}m_{\rm P}^{2}(q-1)q}~~\Rightarrow~~\sigma_{\star}=2^{3(q-2)\over 4(q-1)}\left(\frac{(q-2)\langle{\psi}\rangle^{-2q/(q-2)}}{c_{2q}m_{\rm P}^{2}N_{\star}(q-1)q}\right)^{\frac{1}{4}{(2-q)\over(q-1)}}.$ (51)
Also the mass scale $M$ reads
$M=2m_{\rm P}^{3/2}\sqrt{\pi qc_{2q}\sqrt{3A_{\rm s}}}\,\sigma_{\star}^{-1/2}\left(\langle{\psi}\rangle/\sigma_{\star}\right)^{q/(q-2)},$ (52)
where $\sigma_{\star}$ is given by Eq. (51). The corresponding spectral observables are
$n_{\rm s0}=1-{3q-2\over 2(q-1)}{1\over N_{\star}}~~\mbox{and}~~a_{\rm s0}=-{3q-2\over 2(q-1)}{1\over N_{\star}^{2}}.$ (53)
The results for $n_{\rm s}$ and $a_{\rm s}$ above can be derived by the corresponding ones in Eq. (49) replacing there $p=q/2$. Numerical and more precise values for the quantities above are accumulated in Table 1 for $q=3,5$ and $7$ with similar observational results as those in smFHI-I. Since even values of $q>2$ yield almost identical results with smFHI-I with $p=q/2$, we restrict our discussion to odd values of $q$. In particular, the simplified version of smFHI-II with $\kappa<\sqrt{4\pi}$ is viable only for $q=3$, where $M_{*}$ can be identified with the Planck mass $m_{P}$. In other cases, this identification requires $\kappa>\sqrt{4\pi}$.
5.2 NSUGRA
In the context of NSUGRA, $V_{\rm I}$ is given by the general form of Eq. (34) with the coefficients in Eq. (37a) – (37d) after the imposition of the conditions in Eq. (38). In practice, only $c_{4K}$ from the coefficients above plays a crucial role for the determination of the inflationary observables. For both smFHI-I or smFHI-II the slow-roll parameters read
$\epsilon\simeq\left(\epsilon_{0}^{1/2}+{c_{4K}\sigma^{3}}/{\sqrt{2}m_{\rm P}^{3}}\right)^{2}~~\mbox{and}~~\eta\simeq\eta_{0}+3c_{4K}(\sigma/m_{\rm P})^{2},$ (54)
where $\epsilon_{0}$ and $\eta_{0}$ are given by Eq. (46) or (50) for smFHI-I or smFHI-II respectively. Inserting $\epsilon$ into the leftmost part of Eq. (39) we can compute $N_{\star}$ with result
$N_{\star}\simeq\frac{m_{\rm P}^{2}}{2c_{4K}\sigma_{\star}^{2}}\begin{cases}{}_{2}F_{1}\left(1;\frac{1-p}{3p-2};\frac{2p-1}{3p-2};-\frac{c_{4K}\sigma_{\star}^{2(3p-4)/(p-1)}}{2pc_{2p}m_{\rm P}^{4}\langle{\phi}\rangle^{2p/(p-1)}}\right)&\mbox{for smFHI-I,}\\[11.38109pt] {}_{2}F_{1}\left(1;\frac{q-2}{3q-4};\frac{2(q-1)}{3q-4};-\frac{c_{4K}\sigma_{\star}^{2(3q-4)/(q-2)}}{2qc_{2q}m_{\rm P}^{4}\langle{\psi}\rangle^{2q/(q-2)}}\right)&\mbox{for smFHI-II.}\end{cases}$ (55)
Here $_{2}F_{1}$ is the Gauss hypergeometric function. Due to these complicate forms of $N_{\star}$ we are not able to solve the expressions above w.r.t $\sigma_{\star}$ and pursue our analytical approach. Assuming, however, that $\sigma_{\star}$ do not deviate a lot from their values in Eq. (47) or (51) for smFHI-I or smFHI-II respectively we can provide analytical approximations that capture the essential behavior of $n_{\rm s}$. Namely we find
$n_{\rm s}\simeq n_{\rm s0}+3c_{4K}\begin{cases}4\left(\langle{\phi}\rangle/{m_{\rm P}}\right)^{2p/(2p-1)}\left(N_{\star}c_{2p}p(2p-1)/(p-1)\right)^{(p-1)/(2p-1)}&\mbox{for smFHI-I,}\\[11.38109pt] 2^{\frac{5q-8}{2(q-1)}}\left(\langle{\psi}\rangle/{m_{\rm P}}\right)^{q/(q-1)}\left(N_{\star}c_{2q}q(q-1)/(q-2)\right)^{(q-2)/2(q-1)}&\mbox{for smFHI-II.}\end{cases}$ (56)
Clearly, we can obtain an elevation of the resulting $n_{\rm s}$ w.r.t $n_{\rm s0}$ provided that $c_{4K}$ in Eq. (37b) is positive. Indeed, taking into account Eq. (38) it is not difficult to show that $c_{4K}>0$ for $|N|<7$.
Figure 1: Values of $n_{\rm s}$ allowed by Eqs. (39), (41) and (44) versus $\alpha$ for smFHI-I, $\beta=-1$ or 2-2 and $p=2,3$ and 44 – the marginalized joint $68\%$ $95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours (left plot). Resulting values of $(-N_{0})$ versus $M$ for the same $\beta$ and $p$ (right plot). Shown is also the applied color coding in both plots for the various $\beta$ and $p$ values in the legend of the right plot.
This insight can be explicitly verified by carrying out a precise numerical analysis taking into account all the coefficients in Eq. (37a) – (37d). We use as input parameters $\alpha$ and $\beta=-1$ or 2-2 for selected values of $p$ (smFHI-I) or $q$ (smFHI-II). Our results are shown in Fig. 1 for smFHI-I and Fig. 2 for smFHI-II. In the left panel of each figure we present the allowed values of $n_{\rm s}$ versus $\alpha$ whereas in the right panels we display the corresponding values of $(-N_{0})$ as a function of the mass scale $M$. The inner and outer shaded regions represent the marginalized joint 68% and 95% c.l. contours from the P-ACT-LB-BK18 data. The choices of $p$ and $q$ are given in the right plots. For both types of smFHI we observe that there is a plethora of $\alpha$ values which lead to P-ACT-LB-BK18-compatible $n_{\rm s}$ for both chosen $\beta$. Their magnitudes are less than $0.3$ and decrease as $|\beta|$ increases. Also $|N_{0}|$ decreases below unity as $|\beta|$ increases. Therefore, low $p,q$ and $|\beta|$ values are better motivated from the point of view of naturality. The relevant mass scale $M$ turns out to be of order $1~{\mbox{\rm YeV}}$. E.g. Fixing $\beta=-1$ we obtain
$2/125<\alpha<477/2000~~\mbox{with}~~0.79<M/{\mbox{\rm YeV}}<0.93$ $\mbox{for smFHI-I\ and $p=2$};$ (57a)
$41/500<\alpha<67/250~~\mbox{with}~~0.65<M/{\mbox{\rm YeV}}<0.79$ $\mbox{for smFHI-II\ and $q=3$}.$ (57b)
In Fig. 3 we also display the $a_{\rm s}$ predictions of smFHI-I (left plot) and smFHI-II (right plot) as a function of $n_{\rm s}$ for the same $\beta$ and $p$ or $q$ values used in Fig. 1 or 2. The consideration of the $n_{\rm s}-a_{\rm s}$ correlations reveals that the negligibly small $|a_{\rm s}|$ values obtained in our setup are consistent not only with its $95\%$ c.l. allowed approximate margin in Eq. (43), but also with the $68\%$ c.l. allowed margin for $0.971\lesssim n_{\rm s}\lesssim 0.976$. Note that similar $a_{\rm s}$ (and $r$) values are encountered within shSUGRA too.
Figure 2: The same as in Fig. 1 but for smFHI-II and $q=3,5$ and 77 (instead of $p$).
Figure 3: Values of $n_{\rm s}$, allowed by Eqs. (39), (41) and (44), versus $a_{\rm s}$ for NSUGRAand smFHI-I [smFHI-II] (left [right] plot) corresponding to the parameter choices discussed in Fig. 1 [Fig. 2]. The marginalized joint 68% [95%] c.l. regions derived from P-ACT-LB data are indicated by the dark [light] shaded contours.
Let us emphasize, finally, that the P-ACT-SPT range in Eq. (4) can be obviously accommodated within NSUGRA too. However, the obtained $\alpha$ and $N_{0}$ values are expected to be somehow suppressed and therefore we consider that this framework is more appropriate to fit the P-ACT-LB-BK18 results in Eq. (3). Nonetheless, we can obtain acceptable solutions even for $\alpha=\beta=0$, i.e., without the presence of $\hat{Z}$ and $^{\hat{K}}$ in Eq. (10) at the cost of a very high $N\sim 800$. Although this value is perfectly acceptable within SUGRA, it has no theoretical support and so it can be considered as less natural.
5.3 Monotonicity of $V_{\rm I}$
An outstanding feature of our proposal is that $V_{\rm I}$ in Eq. (34) remains monotonic during smFHI for both SUGRA settings. To highlight it we display the variation [the derivative w.r.t $\sigma$] of $V_{\rm I}$ as a function of $\sigma$ in the left [right] plot of Fig. 4. We focus on smFHI-I with $p=2$. For shSUGRA (solid lines) we use the parameters shown in the column with $p=2$ in Table 1 whereas for NSUGRA (dashed lines) we use $\alpha=1/15$ and $\beta=-1$ resulting to $N_{0}\simeq-1.76$ and $n_{\rm s}=0.9723$. The values of $\sigma_{\star}$ and $\sigma_{\rm f}$ are also depicted. Note that $\sigma_{\star}$ and $\sigma_{\rm f}$ coincide for the two presented lines and so our assumption for the derivation of Eq. (56) is justified. Most importantly, it is clear that $V_{\rm I}$ remains a monotonically increasing function of $\sigma$ in both cases since $V_{\rm I,\sigma}>0$. So, unnatural restrictions on the initial conditions for smFHI due to the appearance of a maximum and a minimum of $V_{\rm I}$ can be avoided.
Figure 4: The variation of $V_{\rm I}$ in Eq. (34) (left panel) and $V_{\rm I,\sigma}$ (right panel) as functions of $\sigma$ for smFHI-I with $p=2$, shSUGRA $\alpha=0$ and $\beta=-3$ resulting to $n_{\rm s}=0.967$ (solid line) or NSUGRA $\alpha=1/15$ and $\beta=-1$ resulting to $n_{\rm s}=0.972$ (dashed line). The values of $\sigma_{\star}$ and $\sigma_{\rm f}$ are also depicted.
6 Conclusions
We have constructed and analyzed two types (I and II) of smFHI (i.e., smooth F-term hybrid inflation) depending on the representation of the Higgs involved fields – see Eq. (5). A key ingredient of our proposal is that the parameters of the superpotential in Eq. (5) are constrained to values that ensure compatibility with the well-established requirement of gauge coupling unification within the MSSM according toEq. (44). This provides a robust theoretical foundation, anchoring our inflationary scenario in a broader particle-physics framework.
However, when confronted with the latest observational constraints of P-ACT-LB-BK18 and P-ACT-SPT data, the model is found to be incompatible within mSUGRA – see Appendix B. To control SUGRA corrections, we incorporate the impact on the inflationary dynamics from a stabilized modulus. It appears only in the Kähler potential of our set-up through the functions $\hat{Z}$ and $^{\hat{K}}$ in Eq. (11) with parameters $\alpha$ and $\beta$ inspired by superstring and D-brane-motivated models. We also considered two classes of Kähler potentials for the inflaton sector in Eqs. (9) and (10) which define the SUGRA settings, shSUGRA and NSUGRA respectively. In shSUGRA, the combination of the shift symmetry for the inflaton-sector together with the adjustment of $\beta$ leads to predictions that coincide with the SUSY version of smFHI and remain consistent with P-ACT-SPT data. In NSUGRA, the conjunction of the inflaton hyperbolic geometry with the variation of $\alpha$ and $\beta$ assist to tame the $\eta$ problem and achieve full consistency with P-ACT-LB-BK18 data – see Figs. 1 and 2. It is gratifying that our solutions are obtained keeping the inflationary potential monotonic – see Fig. 4.
Let us, finally, note that a complete inflationary scenario must also account for the transition to a radiation-dominated universe and the generation of the observed baryon asymmetry. These aspects have been extensively explored in the context of smFHI with canonical or quasi-canonical Kähler potentials – see, e.g., Refs. [5, 9, 15, 10]. Our setup preserves many of the successful features of this post-inflationary evolution, which may impose additional constraints on the viable parameter space and help identify the most compelling version of smFHI.
Appendix A Stabilization of the Modulus Before smFHI
We here present a sample stabilization of $h$ working along the ideas of Ref. [78, 79]. Namely, we assume that $h$ transforms non-trivially under a (gauged) anomalous $U(1)_{\rm FI}$ symmetry carrying charge $q_{h}$. Therefore, the function $f(h)$ in Eq. (11) must be a linear function of $\ln h$. In particular, we assume
$f(h)=\ln(h/m_{\rm P})~~\mbox{with}~~h>m_{\rm P}.$ (58)
Therefore $\hat{Z}$ and $^{\hat{K}}$ in Eq. (11) take the form
$\hat{Z}=\ln|h/m_{\rm P}|^{2}~~\mbox{and}~~{\hat{K}}=\beta m_{\rm P}^{2}\ln\left(\ln|h/m_{\rm P}|^{2}\right).$ (59)
The contribution to the D-term SUGRA potential due to $U(1)_{\rm FI}$ is written as [79, 78]
$V_{\rm FI}=g^{2}{\rm D}^{2}_{\rm FI}~~~\mbox{with}~~~{\rm D}_{\rm FI}=q_{h}K_{h}h+m_{\rm FI}^{2},$ (60)
where we add a Fayet-Iliopoulos term $m_{\rm FI}^{2}$ which naturally assumes values of order $M_{\rm S}$ with contributions from the non-vanishing $U(1)_{\rm FI}$ charges of the various fields of the complete theory. Substituting $K$ in Eq. (8) into Eq. (60) we obtain
$V_{\rm FI}=g^{2}m_{\rm P}^{4}\left(q_{h}\beta\ln|h/m_{\rm P}|^{-2}+m_{\rm FI}^{2}/m_{\rm P}^{2}\right)^{2},$ (61)
where we take into account that $S\ll h$ and $K_{h}\simeq{\hat{K}}_{h}$. A global Minkowski vacuum occurs for
$m_{\rm FI}^{2}\langle{\ln|h/m_{\rm P}|^{2}}\rangle_{\rm I}=-q_{h}\beta m_{\rm P}^{2}~~~\Rightarrow~~~\langle{h}\rangle_{\rm I}=m_{\rm P}\exp(-q_{h}\beta m_{\rm P}^{2}/2m_{\rm FI}^{2}).$ (62)
It is easy to verify that $\langle{V_{{\rm FI},h}}\rangle_{\rm I}=0$ and $\langle{V_{{\rm FI},hh}}\rangle_{\rm I}=0$. Imposing the normalization condition of Eq. (12), we find
$\frac{q_{h}\beta m_{\rm P}^{2}}{2m_{\rm FI}^{2}}=\frac{1}{2}~~~\Rightarrow~~~m_{\rm FI}=\sqrt{-q_{h}\beta}m_{\rm P}.$ (63)
Given that values $\beta=-1$ and 2-2 are encountered in our scheme, we may achieve $m_{\rm FI}\sim M_{\rm S}$ if $q_{h}\sim 1/10$.
Appendix B GUT-Scale smFHI within mSUGRA
Cases Type I smFHI Type II smFHI
$p=2$ $p=3$ $p=4$ $q=3$ $q=5$ $q=7$
$\sigma_{\star}/10^{17}~{\mbox{\rm GeV}}$ $3.165$ $4.57$ $4.965$ $1.556$ $3.11$ $3.683$
$M_{*}/M_{\rm S}$ $0.85$ $0.17$ $0.11$ $5.4$ $0.09$ $0.05$
$\kappa$ $\sqrt{4\pi}$ $0.9$ $\sqrt{4\pi}$
$M/10^{15}~{\mbox{\rm GeV}}$ $1.9$ $3.2$ $3.6$ $0.85$ $1.9$ $2.4$
$\sigma_{\rm f}/10^{17}~{\mbox{\rm GeV}}$ $1.34$ $1.45$ $1.42$ $0.85$ $1.22$ $1.24$
$N_{\rm I*}$ $49.3$ $49.6$ $49.7$ $48.7$ $49.1$ $49.7$
$n_{\rm s}$ $1.04$ $1.1$ $1.12$ $0.987$ $1.03$ $1.06$
$\alpha_{\rm s}/10^{-3}$ $1.65$ $4.5$ $5.9$ $0.87$ $16$ $2.4$
$r/10^{-5}$ $1.2$ $9.6$ $1.5$ $0.048$ $1.2$ 33
$m_{\rm I}/10^{15}~{\mbox{\rm GeV}}$ $0.37$ $1.5$ $2.6$ $0.07$ $0.59$ $1.3$
Table 2: Input and output parameters compatible with Eqs. (39), (41) and (44) within mSUGRA for smFHI-I or smFHI-II and selected $p$ or $q$ respectively.
To appreciate the improvements offered by our versions of smFHI regarding the inflationary predictions, we compare them with the corresponding ones within mSUGRA [2, 16]. In the latter case, the inflationary part of Kähler potential, $K_{\rm I}$ in Eq. (8) is given by
$K_{\rm I}=|S|^{2}$ (64)
and the coefficients in expansion of Eq. (34) read [2]
$c_{2K}=c_{6K}=c_{8K}=0~~\mbox{and}~~c_{4K}=1/2.$ (65)
We remark that the $\eta$ problem is spontaneously solved since we obtain $c_{2K}=0$ by construction. Our proposals are directly comparable with mSUGRA since the same relation is achieved by imposing the conditions in Eq. (35b) or (38) within shSUGRA or NSUGRA, respectively. On the other hand, $c_{4K}$ turns out to be rather high in mSUGRA. Note that $c_{4K}=0$ within shSUGRA and $0.046$ within NSUGRA for the inputs of Fig. 4.
Our results for mSUGRA are presented in Table 2 for smFHI-I or smFHI-II and the same $p$ or $q$ values adopted in Table 1 and Fig. 1 and 2. We remark that $\sigma_{\star}$ and $M$ slightly increase relatively to their values in shSUGRA whereas $|a_{\rm s}|,r$ and $m_{\rm I}$ result one order of magnitude larger. These ramifications shift upward $n_{\rm s}$ which lies well above the observational upper bounds of Eqs. (3) and (4). Only in the case with $q=3$, where the simplified version of smFHI in Ref. [13, 14] is possible, $n_{\rm s}$ turns out to be close to the upper bound in Eq. (3). Let us emphasize, once more, that the constraint of Eq. (44) plays a crucially role in this analysis. Indeed, relaxing this requirement, we may achieve predictions closer to the SUSY results discussed in Sec. 5.1. This limitation of mSUGRA has also been highlighted in Ref. [16]. As a bottom line, GUT-scale smFHI within mSUGRA can be observationally excluded.